RFBG Re-Commissioning with AGW2: 22 September 2010 UT

Observer: JHill (LBTO)
Telescope Operator: DHuerta (LBTO)
Telescope Support: JLittle (LBTO)
Instrument Support: JMorris (LBTO)

Summary

Lost to weather..........clouds and rain in all directions.

Preparations

Torsten supplied at patch to GCS in AO5 to correct some bugs to be seen by MODS tomorrow.

23:55 Paul had to stop and restart OSS to get proper communication with DX M2 UMAC. UMAC was on and pingable, but commands from OSS were not getting through.

Dave and Mike and Gene repaired the damaged connector on the excitation circuit of the Azimuth strip encoder.

01:xx Power glitch from the thunderstorm shows that workstation obs2 is not plugged into the UPS. The situation has been rectified -- IT 2854.

Discovered that IRTC2 was not mounted on the telescope today. Good thing it is raining. Will coordinate with Armando/Simone for when it should next be mounted on the telescope.

Last night's alignment numbers (Tip=+75) for the SX M3 mirror have been memorialized in the collimation models SXTMLUCIFERCollimation.dat and SXTMIRTCFCollimation.dat where they will show as Global Offsets.

Last night's alignment numbers (Tip=+9 Tilt=-35) for the SX M2 mirror have been absorbed into the Global Offset position for M2. This helps by moving the mirror closer to the center of the hexapod range of motion. The new SX M2 Global Offset position is
X=-6.244 Y=-3.060 Z=-1.000 RX=237.030 RY=329.100
These numbers have been saved in SXSMLUCIFERCollimation.dat and SXSMIRTCFCollimation.dat. There is presently no temperature or elevation dependence for M2.

The pointing zeropoints IE=-81 CA=+20 corresponding to the above M2 position have been applied in LUCIFERSX.20100922.ptmod. (Nothing was changed for the IRTCF pointing model.)

Details

Heavy clouds with scattered rain showers around, so daytime engineering work on MODS1, Az Drive, and AdSec Retroreflector continues into the evening. No chance of observing. Both MODS team and AO team seem to have plenty of tests to keep themselves busy.

-- JohnHill - 21 Sep 2010
Topic attachments
I Attachment Action Size Date Who Comment
LUCIFERSX.20100922.ptmodptmod LUCIFERSX.20100922.ptmod manage 325 bytes 22 Sep 2010 - 03:37 JohnHill Updated LUCI1 - LFBG Pointing Model
SXSMLUCIFERCollimation.20100922.datdat SXSMLUCIFERCollimation.20100922.dat manage 1 K 22 Sep 2010 - 03:40 JohnHill Updated LUCI1 - LFBG M2 Collimation Model
SXTMLUCIFERCollimation.20100922.datdat SXTMLUCIFERCollimation.20100922.dat manage 880 bytes 22 Sep 2010 - 03:39 JohnHill Updated LUCI1 - LFBG M3 Collimation Model
Topic revision: r4 - 22 Sep 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback